Hi SPARO/ACBAR collaborators:

Hua-bai, Bob L., Paolo, and I have been working at Pole for about two
weeks now.  Bill, Mike, and Martin left a few days ago.  Mark was here at
Pole for a few days, but had to return to US due to a family emergency.  
He will return for the last three weeks of the season.  Matt N. and Bob P.
will be here for the last two weeks.

Bill et al. left us a telescope in good working order.  Martin's new and
improved electronics heater system has been running (in analog mode) with
no problems for over a week.  We haven't been running the mirror heating
system lately, though it was tested earlier.  When Mark returns we'll
re-test their computer mode for the electronics heater system and also
re-test their mirror heating system.  The chopper has been working almost
flawlessly since I've been here.  Mike and Bill now have two optical
encoders on it, and a spare ready. Bill installed the third AZ Techron and
tuned up the AZ servo control.  He improved the stability by making
changes to the PID circuit.  Specifically, he found and replaced a bad
chip, changed a resistor in the I circuit, and turned the I and D terms
down to zero.

ACBAR came off five days ago. SPARO is in the lab, at 300 mK.  Cryostat
working well.  Bob L. has successfully tested four out of five of the
computer interfaces we need.  Greg G. is providing full-time stateside
support for SPARO software.  It was helpful that he and Megan got a
"state-side simulator" working at NU in Fall.  The only computer interface
that remains to be tested is the interface to the frame grabber.

Using Hua-bai's new template, we have been reconfiguring the telescope
optics to correspond to the SPARO configuration.  we discovered that i
made a sign error back in Fall when I told Bill what chopper translation
we would need to get from the ACBAR to the SPARO configuration.  As a
result, the chopper position needed for SPARO was out of range.  Luckily,
the fix was very easy.  Hua-bai remade his SPARO template in such a way as
to allow us to operate the chopper at almost exactly the same
translational postion as for ACBAR.  There is no lost accuracy and no
down-side to this solution.  We should be done with optics tomorrow.

Paolo and Allan Day have been trained to install ACBAR, and Paolo trained
on ACBAR observations.

We could be installing SPARO in a few days, though i may want to keep it
in the lab for a while longer to experiment with controlling the flow
through the capillary.  SPARO has always had the problem that the
capillary flow can vary from very high (2 milliTorr at the pump input) to
very low (0.05 Torr at pump input).  If the flow is higher than about 0.5
Torr, we get bad microphonics, but if it is lower than about 0.1 Torr, we
can't do the He-3 cycle.  If the flow is too low, we can raise it by
letting SPARO warm to 90 K while it is on the telescope.  In past, there
has never been a problem with too high flow on the telescope, as the flow
eventually settles withiin the desired range.  but we're always concerned
that it could settle at a value that is too high.  Right now, for example,
it has been at 2 milliTorr and stable for a couple of days.  So i may
spend a few more days playing with the flow in the lab to see if i can
learn how to control it better.

All in all, things are going very well.